OpenStudy (anonymous):

find the vertex f(x)=2x^2+2x+5

OpenStudy (anonymous):

Formula =-b/2a a=2 b=2 =-1/2

OpenStudy (anonymous):

how about for y

OpenStudy (anonymous):

9/2